A mixture of oxygen and hydrogen is analyzed by passing it over hot copper oxide and through a drying tube. Hydrogen reduces the CuO according to the equation, CuO + H2 Ò Cu + H20 ; Oxygen then oxidizes the copper formed : Cu + ½O2 Ò CuO. 100 cm³ of the mixture measured at 25°C and 750 mm yields 84.5 cm³of dry oxygen measured at 25°C and 750 mm after passing over CuO and drying agent. What is the mole present of H2 in the mixture?

Grade:11

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi nikit,

V=100 ml

let H2 is x ml

then xml of H2 and x/2 ml O2 are consumed.....

then (100-x) ml O2 initially

and 100-x-x/2 ml O2 finally

therefore,

100-x-x/2 = 84.5

3x/2 = 15.5

x=31/3

mole% of H2 in the mixture is (31/3)*100/100 = 10.33%